# Depth to Anatomy: Learning Internal Organ Locations from Surface Depth Images

### v1

@@ 3.1 Data Preparation
spleen scapula left vertebrae_T12
kidney right scapula right vertebrae_T11
kidney left clavicula left vertebrae_T10
liver clavicula right vertebrae_T9
stomach femur left vertebrae_T8
pancreas femur right vertebrae_T7
lung right hip left vertebrae_T6
lung left hip right vertebrae_T5
trachea sacrum vertebrae_T4
thyroid gland vertebrae L5 vertebrae_T3
duodenum vertebrae L4 vertebrae_T2
urinary bladder vertebrae L3 vertebrae_T1
@@ 4 Experiments and Results
Figure 4: Dice coefficient for volumetric organs segmentation. Values are averaged for bilateral organs (e.g., kidneys, lungs) and vertebral subgroups (thoracic and lumbar vertebrae). Pix2Vox significantly outperforms the baseline, demonstrating its ability to generate accurate, patient-specific organ segmentations from a single coronal depth image.
@@ 4.1 Training and evaluation details
Figure 5: Average Symmetric Surface Distance (ASSD) for volumetric organs segmentation. Values are averaged for bilateral organs (e.g., kidneys, lungs) and vertebral subgroups (thoracic and lumbar vertebrae). Pix2Vox exhibits markedly lower surface distances compared to the mean model, indicating superior boundary accuracy and shape fidelity.
@@ 4.2 Baselines
Figure 6: Qualitative comparison between predicted and ground truth volumetric segmentations. The Pix2Vox model demonstrates a strong ability to accurately infer both the shape and location of internal organs solely from a single depth image. The examples highlight successful predictions for both bony (vertebrae and scapula) and soft tissue organs (lung and heart) across different organ sizes and positions.
@@ 4.3 Metrics
We evaluate the localization and segmentation performance of all models using a combination of complementary metrics that capture both spatial accuracy and anatomical shape fidelity. Bounding boxes are derived from the volumetric segmentation masks produced by the mean model, the Pix2Vox model, and the three-plane predictions of the 2.5D CNN. To quantify localization accuracy, we calculate the Mean Absolute Detection Offset Error (DOE), defined as the absolute distance between the corresponding faces of the predicted and ground truth bounding boxes along each axis. To better understand model behavior in different spatial dimensions, DOE is reported separately for each side of the bounding box (e.g., left, right, anterior, posterior, superior, inferior). This directional breakdown allows us to assess how accurately each method positions the anatomical structures along the full 3D extent. Importantly, this metric reflects practical constraints in clinical imaging scenarios, particularly in automated patient positioning and scan planning, where under- or over-estimating organ boundaries can lead to suboptimal coverage or increased scan time. To assess the model’s ability to reconstruct organ shape, we compute the Average Symmetric Surface Distance (ASSD) between the predicted and ground truth segmentations, which captures discrepancies at the organ boundary. In addition, we report the Dice coefficient as a standard volumetric overlap metric. Both Dice and ASSD are calculated only for models that produce full 3D volumetric predictions, enabling us to evaluate how well the organ shape can be recovered from 2D depth image inputs alone.
@@ 4.4 Results
Quantitative analysis using the Detection Offset Error (DOE) highlights the strong performance of the proposed Pix2Vox model in localizing internal organs in 3D space based solely on a single coronal depth image (Figure 3). Specifically, in the coronal plane, Pix2Vox demonstrates a more than twofold improvement in mean absolute DOE compared to the mean model baseline, achieving an average localization error of less than 10mm across all evaluated organs. However, performance degrades when localizing organs along the anterior-posterior (depth) axis, where the model shows higher variance and larger prediction errors. This limitation is not unexpected, given that the coronal depth image inherently provides less depth-specific information. Nonetheless, the ability of the CNN to infer organ depth from contextual and structural cues in a single 2D image remains noteworthy, demonstrating the model’s strong capacity for generalization. The 2.5D CNN approach, which integrates predictions from coronal, sagittal, and axial projections, exhibits superior localization performance in the coronal plane and achieves more robust bounding box estimation in lateral and vertical dimensions. However, it still underperforms in predicting organ position along the depth axis. This shortfall may be attributed to the limited representational capacity of the 2D projection-based architecture and the absence of a fully 3D decoding mechanism, which constrains the model’s ability to jointly reason across spatial dimensions.
Quantitative evaluation using the Dice coefficient and ASSD)(Figure 4 and Figure 5), alongside qualitative visualizations of Pix2Vox predictions (Figure 6), reveals the model’s remarkable ability to infer internal organ shape from a single 2D depth image. Unlike the mean model baseline, which simply reflects anatomical averages, Pix2Vox demonstrates a substantial improvement in both overlap and surface accuracy metrics. This performance gap highlights the model’s capacity to move beyond population-level anatomical assumptions and instead learn spatially-aware, patient-specific volumetric representations directly from depth cues present in surface imagery. The model’s strong performance across diverse organ types, sizes, and locations further supports the idea that it has learned robust spatial priors, enabling generalization across a heterogeneous population.
These findings suggest that internal organ anatomy can be successfully estimated directly from body surface depth information, highlighting its value as a non-invasive surrogate for internal imaging. This demonstrates the potential of integrating camera-based organ localization into radiological workflows, where such technology could enhance scan planning and optimize patient positioning.
@@ 5 Discussion
In this study, we introduce an approach for automated patient positioning that leverages body surface depth images to estimate the 3D position and shape of internal organs. Our quantitative and qualitative results demonstrate the feasibility of using a single depth image to approximate the internal anatomical layout. The proposed Pix2Vox model successfully learns to extract cues from surface morphology, such as torso shape, curvature, and contour depth, to predict the volumetric extent and spatial orientation of internal organs.
The model’s strong localization accuracy across multiple organs a promising direction to support scan planning and optimize table positioning without the need for external markers or scout imaging. Automated positioning could reduce reliance on trial-and-error adjustments by radiology technicians, potentially decreasing scan time, improving reproducibility, and reducing inter-operator variability. Additionally, the model’s capacity to generalize across diverse patient anatomies, as evidenced by its performance on a large and heterogeneous test set, underlines its robustness in dealing with physiological variation.
Despite these encouraging findings, several limitations and opportunities for future research must be acknowledged:
@@
  • Domain gap between synthetic and real-world depth data: The depth images used in this study were synthetically generated from MRI-derived segmentations and surface reconstructions, which may not fully capture the variability present in real-world depth camera outputs. In clinical environments, additional sources of noise, such as occlusions, limited spatial resolution, lighting artifacts, and sensor-specific distortions, can affect image quality. Furthermore, patients may be covered with blankets, wear hospital gowns, or use assistive devices, all of which can obscure key body surface features. Bridging this domain gap is crucial for real-world applicability and may require domain adaptation techniques or model fine-tuning using real sensor-acquired data.
@@
  • Assumption of consistent patient pose and orientation: The training data used in this work consisted of patients in a standardized supine position with minimal variation in posture. Although position-based data augmentations were employed during training to increase robustness, real-world scenarios often involve a wider range of patient poses due to comfort needs or mobility restrictions. The current model may be sensitive to these variations, and further research is needed to evaluate and improve its generalizability to diverse body orientations and clinical settings.
@@
    The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.

### v2

@@ 2.4 Processing of Real-World Depth Images
spleen scapula left vertebrae T12
kidney right scapula right vertebrae T11
kidney left clavicula left vertebrae T10
liver clavicula right vertebrae T9
stomach femur left vertebrae T8
pancreas femur right vertebrae T7
lung right hip left vertebrae T6
lung left hip right vertebrae T5
trachea sacrum vertebrae T4
thyroid gland vertebrae L5 vertebrae T3
duodenum vertebrae L4 vertebrae T2
urinary bladder vertebrae L3 vertebrae T1
@@ 3 Experiments and Results
Figure 4: Dice coefficient for volumetric organs segmentation. Values are averaged for bilateral organs (e.g., kidneys, lungs) and vertebral subgroups (thoracic and lumbar vertebrae).
@@ 3.1 Training and evaluation details
Figure 5: Average Symmetric Surface Distance (ASSD) for volumetric organs segmentation. Values are averaged for bilateral organs (e.g., kidneys, lungs) and vertebral subgroups (thoracic and lumbar vertebrae).
@@ 3.2 Baselines
Figure 6: Qualitative comparison between predicted and ground truth volumetric segmentations.
@@ 3.3 Metrics
We evaluate the localization and segmentation performance of all models using a combination of complementary metrics that capture both spatial accuracy and anatomical shape fidelity. Bounding boxes are derived from the volumetric segmentation masks produced by the mean model, the Pix2Vox model, and the three-plane predictions of the 2.5D CNN. To quantify localization accuracy, we calculate the Mean Absolute Detection Offset Error (DOE), defined as the absolute distance between the corresponding faces of the predicted and ground truth bounding boxes along each axis:
$\text{DOE}=\frac{1}{N}\sum_{i=1}^{N}|\text{face}_{i,\text{pred}}-\text{face}_{i,\text{gt}}|$ (6)
where $\text{face}\in\{x_{min},x_{max},y_{min},y_{max},z_{min},z_{max}\}$ represents the coordinates of the six bounding box planes. To better understand model behavior in different spatial dimensions, DOE is reported separately for each side of the bounding box (e.g., left, right, anterior, posterior, superior, inferior). This directional breakdown allows us to assess how accurately each method positions the anatomical structures along the full 3D extent. Importantly, this metric reflects practical constraints in clinical imaging scenarios, particularly in automated patient positioning and scan planning, where under- or over-estimating organ boundaries can lead to suboptimal coverage or increased scan time. To assess the model’s ability to reconstruct organ shape, we compute the Average Symmetric Surface Distance (ASSD) between the predicted and ground truth segmentations, which captures discrepancies at the organ boundary. In addition, we report the Dice coefficient as a standard volumetric overlap metric. Both Dice and ASSD are calculated only for models that produce full 3D volumetric predictions, enabling us to evaluate how well the organ shape can be recovered from 2D depth image inputs alone.
@@ 3.4 Results
3.4.1 Localization Performance
Quantitative analysis using DOE highlights the balanced 3D localization performance of the proposed Pix2Vox model across all anatomical axes, as illustrated in Figure 3. Within the coronal plane, Pix2Vox demonstrates a substantial improvement over the population-based mean model. Specifically, it achieves a mean absolute error of $7.87\pm 7.69$ mm for the right boundary and $6.87\pm 6.62$ mm for the left, significantly outperforming the baseline values of $19.86\pm 17.14$ mm and $21.23\pm 17.19$ mm, respectively.
The 2.5D CNN approach, leveraging multi-planar projections, achieves the highest precision in lateral localization with a left-side DOE of $2.79\pm 2.93$ mm. While both deep learning approaches exhibit comparable performance along the superior-inferior axis, with errors ranging between $8.39$ mm and $11.36$ mm, the anterior-posterior axis remains the most challenging dimension for the 2.5D method. In the anterior direction, Pix2Vox yields an error of $13.29\pm 8.41$ mm, which is notably superior to the $21.31\pm 9.53$ mm error produced by the 2.5D CNN. Interestingly, the mean model provides a competitive baseline along the depth axis, yielding errors of $13.00\pm 10.63$ mm in the anterior direction and $12.90\pm 12.48$ mm in the posterior direction.
Figure 7: Qualitative evaluation on real-world depth images of volunteers positioned in an MRI scanner, following the setup described in Section 2. The predicted internal structures are displayed as volumetric segmentations projected onto the coronal plane. From left to right: (1) input depth image; (2) overlay of lungs, kidneys, and urinary bladder; (3) overlay of trachea, heart, spleen, liver, and pancreas; (4) overlay of thyroid gland, aorta, stomach, and duodenum; and (5) volumetric segmentation of the skeletal system, shown in a different perspective for each subject.
3.4.2 Segmentation and Shape Fidelity
The volumetric reconstruction performance was further evaluated using the Dice similarity coefficient and ASSD. The Pix2Vox model achieved an overall mean Dice score of $0.48\pm 0.20$, representing a clear advancement over the population-based mean model baseline of $0.34\pm 0.21$. Anatomical analysis reveals that large organs with a high correlation to the body surface exhibit the most robust reconstruction. Specifically, the lungs and liver achieved Dice scores of $0.78\pm 0.04$ and $0.71\pm 0.07$, respectively. In contrast, smaller or highly variable structures, such as the thyroid gland at $0.24\pm 0.15$ and the trachea at $0.32\pm 0.19$, presented greater challenges for surface-based volumetric estimation. Skeletal structures also showed competitive performance, with the femur at $0.62\pm 0.10$ and the sacrum at $0.61\pm 0.15$ yielding stable overlap metrics.
The surface fidelity metrics follow a similar trend. Pix2Vox reduced the overall mean ASSD to $8.34\pm 8.11$ mm, compared to $13.17\pm 10.84$ mm for the mean model. The highest surface accuracy was observed for the scapula and the hip, with errors of $5.96\pm 1.65$ mm and $6.08\pm 1.27$ mm, respectively. Notably, while some internal organs exhibit lower Dice scores due to their small volume, the corresponding ASSD values remain relatively low, as seen with the $6.58$ mm error for the thyroid gland.
3.4.3 Evaluation on Real-World Depth Images
To assess the translational potential of our framework, we evaluated the trained model on depth maps acquired from volunteers using a Microsoft Kinect v2 sensor. As illustrated in Figure 7, the model reconstructs internal anatomical structures that appear spatially consistent with the subject’s external surface manifold. The predicted organ locations and global orientations align with expected human anatomy, suggesting that the model is robust to the real-world depth images.
However, detailed visual inspection reveals specific reconstruction artifacts inherent to surface-only volumetric estimation. In the skeletal overlays, particularly in the middle image of Figure 7, morphological artifacts are noticeable in the femoral regions, appearing as unusually narrowing towards bottom of the image. Furthermore, the lung segmentations in the middle and bottom rows exhibit localized peak-like artifacts. While the overall abdominal arrangement is preserved, the morphology of the liver in middle image shows a loss of characteristic lobar sharpness.
@@ 4 Discussion
The quantitative and qualitative results indicate that 3D internal anatomy can be estimated from 2D surface depth information, though the degree of accuracy is contingent upon the specific organ of interest and the anatomical axis.
Analysis of the Detection Offset Error reveals distinct trade-offs between the evaluated architectures. While the Pix2Vox model provided more balanced 3D localization across all axes, the 2.5D CNN achieved the highest precision in the lateral left-right dimensions. This is likely attributable to its multi-planar projection approach, which excels at identifying boundary edges in 2D space. However, the performance degradation of the 2.5D CNN along the anterior-posterior axis suggests that the absence of a dedicated 3D decoding mechanism limits its ability to resolve depth-specific structural cues. The lower performance in the anterior-posterior direction observed for both deep learning approaches is not unexpected, as a single-view coronal depth map provides limited information regarding the posterior extent of internal organs.
The disparity in Dice scores between large organs, such as the lungs, and smaller structures, such as the thyroid gland, suggests that volumetric overlap is highly dependent on the absolute size of the target. However, the relatively low ASSD for small organs indicates that the model accurately localizes the anatomical center of these structures even when it cannot precisely resolve their morphological boundaries.
4.1 Real-World Generalization and Artifacts
The evaluation on real-world Kinect v2 data (Section 3.4.3) demonstrates that the model is robust to real-world depth acquisition, but it also highlights inherent limitations in surface-to-volume inference. The narrowing artifacts observed in the femoral regions and the peak-like artifacts in the lung segmentations suggest a sensitivity to surface curvature and sensor noise. The loss of lobar sharpness in the liver morphology further indicates that internal soft-tissue boundaries are not fully constrained by the surface manifold. These discrepancies suggest that while global positioning is effectively captured, fine-grained boundary reconstruction of non-surface-correlated tissues remains a challenge.
4.2 Limitations and Clinical Considerations
Several technical constraints must be addressed for clinical implementation:
@@ 4.2 Limitations and Clinical Considerations
  • Domain Gap: The depth images used in this study were synthetically generated from MRI-derived segmentations and surface reconstructions, which may not fully capture the variability present in real-world depth camera outputs. In clinical environments, additional sources of noise, such as occlusions, limited spatial resolution, and sensor-specific distortions, can affect image quality. Furthermore, patients may be covered with thick blankets or use assistive devices that obscure key body surface features. Bridging this domain gap is crucial for real-world applicability and may require domain adaptation techniques or model fine-tuning using real sensor-acquired data.
@@
  • Positional Constraints: The training data used in this work consisted of patients in a standardized supine position with minimal variation in posture. Although geometric data augmentations were employed during training to increase robustness, real-world scenarios often involve a wider range of patient poses due to comfort needs or mobility restrictions. The current model may be sensitive to these variations, and further research is needed to improve its generalizability to diverse body orientations.
In conclusion, the results demonstrate the technical feasibility of using depth-based volumetric estimation for automated isocenter alignment. While the precision is promising for the automated table positioning use-case, the observed morphological artifacts suggest that surface data alone may be insufficient for high-fidelity reconstruction of all anatomical structures.
@@
The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of human subjects.
@@
Appendix A Detailed Performance Metrics
A detailed breakdown of performance metrics for each organ. Figures 1013 present the evaluation across all 41 anatomical structures.
Figure 8: Dice similarity coefficient for internal organs.
Figure 9: Dice similarity coefficient for internal organs.
Figure 10: Surface distance metrics for internal organs.
Figure 11: Surface distance metrics for internal organs.
Figure 12: Detection offset for internal organs.
Figure 13: Detection offset for internal organs.